OpenStudy (anonymous):

The local minimum of the cubic function: y= -1/3x^3 + 6x^2 - 27x - 9 is found when x=?

OpenStudy (anonymous):

take the derivative, set it equal to zero and solve

OpenStudy (anonymous):

this is a cubic with negative leading coefficient, so it looks like 

OpenStudy (anonymous):

local min therefore is located when \(x\) is the smaller of the two zeros. the larger of the two will be give the local max

OpenStudy (anonymous):

x=-2.25

OpenStudy (anonymous):

-x^2+12x-27 = 0 -x^2+12x=27 -x^2+x=27/-12 x=-2.25

OpenStudy (anonymous):

right?